# Chance Constrained PDE-Constrained Optimal Design Strategies Under High-Dimensional Uncertainty

### v1

@@ 5 Numerical Results
This section presents numerical experiments utilizing the proposed design under uncertainty framework for beam-insulation scenario in Figure 1 in both 2D and 3D. The objective is to achieve thermal insulation and mechanical stability avoiding stress concentration while mitigating uncertainty in the design process. The analyses presented in this section, investigate the accuracy and convergence of quadratic and Monte Carlo approximation of design objective and explores the effect of different limiting critical stress, critical chance, and uncertain parameter variance and spatial correlation on the optimal design, spatial pattern of $d_{\text{opt}}$, as well as the probability distributions of the design objectives $Q$. The parameters of the forward model are summarized in Table 1. The computational implementation of the framework relies on several open-source libraries like FEniCS [56] for the finite element solution of the forward model, SOUPy [57] for the quadratic approximation of the design objective and implementation of chance constraint, and hIPPYLib [54] for the Gaussian random field, trace estimator and INCG optimizer.
@@ 5.1 Spatially-correlated uncertain parameter
Refer to caption  Refer to caption  Refer to caption  Refer to caption
Refer to caption  Refer to caption  Refer to caption  Refer to caption
[…]
(a)                                              (b)
Refer to caption            Refer to caption
(c)                                              (d)
Figure 5: The effect of boundary on the Gaussian random field in (11) for the scenario used in Figure 7: (a) marginal variance of the anisotropic uncertain parameter $m$, and point-wise correlation structure centered at (b) point (0.5,0.75) (c) point (0.25,0.75) (d) point (0.25,0.25). Coordinates $(x,y)$ are defined with respect to the bottom left corner of the beam-insulator domaim. The uncertain parameter $m$ is characterized by a mean value of $\bar{m}=0$, a variance of $\sigma^{2}=0.5^{2}$ and a correlation length of $L_{c}=0.25$.
@@ 5.2 Quadratic and Monte Carlo approximations
To assess the accuracy and efficiency of the quadratic approximation and Monte Carlo estimations of mean and variance of design objective with a control variate on the optimal design solution, we perform comparative analyses. As the sample size and number of dominant eigenvalues increase, both methods converge to similar mean and variance values, consistent with standard Monte Carlo estimations, as shown in Figure 6. For large sample sizes, the estimated mean and variance from the standard Monte Carlo method (Figure 6(a)) and the Monte Carlo method with control variate (Figure 6(b)) are nearly identical. However, the quadratic approximation as a control variate achieves lower error bounds with significantly fewer samples, as evident in the box plots. Furthermore, with only $N_{tr}=25$ trace estimators, the quadratic approximation closely matches the Monte Carlo results (Figure 6(c)), leading to significant savings in computational cost without loss in accuracy.
@@
Figure 6: Convergence plots corresponding to scenario in Figure 7 for the mean and variance of the design objective $Q$ : (a) standard Monte Carlo method (b) Monte Carlo with control variate and (c) Quadratic approximation. The uncertain parameter $m$ has the correlation length $L_{c}=0.05$ and a variance of $\sigma^{2}={0.25}^{2}$.
Figures 7 and 8 compare the effects of different variances and correlation lengths of the uncertain parameter $m$ on the optimal design $d_{\text{opt}}$ indicating the spatial distribution of aerogel porosity and the probability distributions of the design objective $Q$ representing the thermal insulation performance of the thermal break. The optimal design is evaluated using standard Monte Carlo, Monte Carlo with control variate, and a quadratic Taylor approximation of the mean and variance of $Q$. In Figure 7 (low variance and short correlation length), the morphologies of $d_{\text{opt}}$ and probability distributions of $Q$ are nearly identical across all three methods. However, in Figure 8 (high variance and long correlation length), the quadratic approximation predicts different optimal design pattern with higher $Q$, while Monte Carlo with control variate yields results closely matching standard Monte Carlo but with significantly fewer samples (PDF solved).
@@
Figure 7: The top row shows the optimal design obtained for $\sigma=0.25$ for the three different approaches: (a) Sampled Average approximation (500 samples) (b) Monte Carlo correction (100 samples) and (c) Quadratic approximation ($N_{tr}=25)$. The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above. The uncertain parameter $m$ has a correlation length $L_{c}=0.05$ for all three cases.
Interestingly, while the spatial pattern of the optimal design in Figure 7 appears nearly symmetric, the higher uncertainty in Figure 8 results in optimal designs that consistently position stronger material (lower aerogel porosity) along the horizontal axis, across all three approximation methods. This pattern reflects the anisotropy of the uncertain parameter, modeled with a higher correlation length along the layer deposition direction (see Figure 4). Also, the optimal designs at mean $\tilde{m}=0$ are much similar in Figure 7 as compared to Figure 8 where there is a deviation in the design patterns, which shows that there is a trade-off favoring the reduction of the uncertainty over mean performance for our design formulation (12). These results demonstrate that the proposed optimization under uncertainty framework effectively captures spatial uncertainty, producing more reliable design solutions.
Figure 8: The top row shows the optimal design obtained for $\sigma=0.5$ for the three different approaches: (a) standard Monte Carlo using 500 samples, (b) Monte Carlo with control variate using 100 samples, and (c) Quadratic approximation using $N_{tr}=25$. The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above. The uncertain parameter $m$ has a correlation length of $L_{c}=0.25$ for all three cases.
@@ 5.3 Scalability with respect to number of design parameters
We assess the scalability of the algorithm for PDE-constrained optimization under uncertainty with respect to the design dimension, determined by the number of nodes in the spatial discretization of $d$, and governed by two primary factors: Firstly, the eigenvalue decay rate of the covariance-preconditioned Hessian, estimated via the randomized trace estimator, and secondly the convergence behavior of the INCG optimizer. Figure 9 illustrates the behavior of the chance constraint function in (14) with the number of steps in the continuation scheme, the gradient norm 64 and the absolute values of the generalized eigenvalues 48 for the last step of chance convergence discussed in Algorithm 3. These metrics are presented for two design parameter dimensions, corresponding to the optimal design problem depicted in Figure 7 (c).
@@
These results demonstrate that the quadratic approximation enables rapid decay of the critical chance, with the gradient norm converging below the tolerance threshold within 100 iterations for both mesh resolutions. The minimal dependence of convergence on mesh resolution highlights that the gradient-based INCG optimizer is effectively independent of the number of uncertain parameters. Additionally, Figure 9(c) shows a rapid eigenvalue decay, unaffected by the dimension of uncertain parameters, supporting the scalability of the randomized eigensolver used for trace approximation.
1st                     2nd                     10th                     20th
Refer to caption  Refer to caption  Refer to caption  Refer to caption
40th                   60th                   80th                   100th
Figure 10: Eigenfunctions of the covariance-preconditioned Hessian of the design objective corresponding to eigenvalue decay for the scenario in Figure 7 (c).
Figure 10 depicts the eigenfunctions of the covariance-preconditioned Hessian for the design objective. Eigenvectors associated with larger eigenvalues capture broader spatial scales of the design parameter, while those linked to smaller eigenvalues become increasingly oscillatory and less significant. Together, these dimension-independent factors suggest that leveraging quadratic Taylor approximations and stochastic optimization techniques enables a scalable approach for PDE-constrained optimal design under uncertainty with the enforcement of chance constraints, regardless of the dimensionality of the uncertain parameters.
@@ 5.4 Optimal design: effect of chance constraint
Figure 11: The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of critical chance: (a) $\alpha_{c}=0.05$ (b) $\alpha_{c}=0.08$ and (c) $\alpha_{c}=0.1$. The bottom row represents the von Mises stress for the corresponding optimal designs in the top row. In all results $T_{cr}=22.5\>\text{MPa},L_{c}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{R}=1\times 10^{-5}$.
@@
Figure 12: The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of limiting critical stress: (a) $T_{cr}=20\>\text{MPa}$ (b) $T_{cr}=22.5\>\text{MPa}$ and (c) $T_{cr}=25\>\text{MPa}$. The bottom row represents the von Mises stress for the corresponding optimal designs in the top row. In all results $\alpha_{c}=0.05,L_{c}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{R}=1\times 10^{-5}$.
The chance constraint function defined in (14) incorporates parameters such as $\alpha_{c}$ and $T_{cr}$, which govern the enforcement of constrain to mitigate stress concentrations within the domain. In particular, the probability of chance $P(f(m,d)\geq 0)$ is limited by a threshold $\alpha_{c}$, where the function $f$ is the difference between the limiting critical stress $T_{cr}$ and the p-norm of the von Mises stress $T_{\text{pn}}$. To examine the influence of $\alpha_{c}$ on optimal design, a parametric study was conducted with constant parameters, as shown in Figure 11. The critical stress limit was set at $T_{cr}=22.5\>\text{MPa}$, while $\alpha_{c}$ was varied among 0.05, 0.08, and 0.1. This figure presents the optimal design, and von Mises stress distribution evaluated at the mean of the uncertain parameter $\tilde{m}=0$. The results indicate that increasing critical chance $\alpha_{c}$ expands the regions of mechanically stronger material (highlighted in red), effectively reducing localized stresses across the component. However, this improvement in mechanical stability comes at the expense of insulation performance of the thermal break.
@@
The limiting critical stress, $T_{cr}$, defines the maximum permissible von Mises stress within the beam structure. Figure 12 demonstrates the influence of $T_{cr}$ on the optimal designs, evaluated at the mean of the uncertain parameter $\tilde{m}=0$, along with the corresponding von Mises stress distributions. The findings indicate that increasing the critical stress $T_{cr}$ results in the allocation of stronger insulating material (depicted in blue) across a larger region of the domain. This enhances the insulation performance but compromises the mechanical strength of the component. Thus, a higher value of $T_{cr}$ improves insulation efficiency at the expense of reduced mechanical stability. Figure 13 shows the state parameters and porosity field evaluated at the mean of the uncertain parameter $\tilde{m}=0$ and three samples for the optimal design in Figure 12 (b). The states include the solid displacement magnitude $||\mathbf{u}_{s}||$, magnitude of effective solid stress $||\mathbf{T}_{s}||$, , solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$ and fluid pressure $p$.
@@ 5.5 Three-dimensional design of beam-insulator system
Figure 14: Convergence of critical chance $\alpha_{c}$ with the results obtained for optimal design at mean $\tilde{m}=0$ for steps in adaptive optimization discussed in Algorithm 3 from step 1 to step 5 shown in (b)-(f). The uncertain parameter $m$ has a correlation length of $L_{c}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$,the critical chance is $\alpha_{c}=0.1$, and the critical stress considered is $T_{cr}=15\>\text{MPa}$.
We demonstrate the efficiency and scalability of the proposed optimization under uncertainty framework by applying it to a three-dimensional design problem. Specifically, we consider the 3D thermal break scenario shown in Figure 1, with a thickness of 0.5. In this case, a uniform circular load is also applied to the middle of the top surface, simulating the influence of additional building envelope assemblies. The uncertain parameter $m$ is modeled with a horizontal correlation length of $L_{c}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$. For the chance constraint, we set a critical chance of $\alpha_{c}=0.1$ and a critical stress of $T_{cr}=15\>\text{MPa}$. The finite element discretization of the 3D domain resulted in 389,044 design parameters.
@@
Figure 15: Final design obtained after five steps of the continuation scheme, achieving convergence of the critical chance $\alpha_{c}$. The 3D domain’s finite element discretization, resulted in 389044 design parameters.
Figure 14 illustrates the optimal designs obtained at each step of the continuation scheme outlined in Algorithm 3. Regions shown in red correspond to material with low porosity ($\phi_{f}=0.1$). As the continuation scheme progresses, the red regions gradually decrease, reducing mechanical support in the domain and increasing the p-norm of von Mises stress. Consequently, the probability in (14) approaches the critical chance threshold of $\alpha_{c}=0.1$. By step 5, the final design, shown in Figure 15, emerges, featuring a column-like structure of low-porosity material $\phi_{f}=0.1$, providing support aligned with the direction of the applied external load.
@@ 6 Conclusions
This study introduces an efficient and scalable computational framework for chance-constrained optimal design under high-dimensional uncertainty governed by partial differential equations (PDEs). The framework optimizes the spatial distribution of a spatially correlated and uncertain porosity field in additively manufactured aerogel thermal breaks to maximize thermal insulation performance while mitigating stress concentrations. Two key strategies are implemented within the framework. The first utilizes a quadratic Taylor expansion of the design objective with respect to the uncertain parameters to efficiently estimate the mean and variance in the cost function, as well as the chance constraints. The second employs the quadratic approximation as a control variate to enhance the efficiency of Monte Carlo estimations. The framework incorporates a randomized algorithm to solve the generalized eigenvalue problem, enabling efficient computation of the trace of the covariance-preconditioned Hessian. Scalability with respect of high-dimensional uncertain parameter is achieved through the rapid decay of eigenvalues and the dimension-independent convergence of the optimization iterations.
@@
This study addresses the computational challenges of PDE-constrained optimization under high-dimensional uncertainty by leveraging intrinsic dimensionality and incorporating chance constraints to prevent stress concentrations. The framework enhances optimal design reliability, offering a robust and scalable approach while paving the way for future advancements to refine and extend its capabilities.

### v2

@@ 5 Numerical Results
This section presents numerical experiments using the proposed framework for design under uncertainty for beam-insulation scenario in Figure 1 in both 2D and 3D. The objective is to achieve thermal insulation and mechanical stability avoiding stress concentration while mitigating uncertainty in the design process. The analyses presented in this section, investigate the accuracy and convergence of quadratic and Monte Carlo approximation of design objective and explores the effect of different limiting critical stress, critical chance, and uncertain parameter variance and spatial correlation on the optimal design, spatial pattern of $d_{\text{opt}}$, as well as the probability distributions of the design objectives $Q$. The parameters of the thermomechanical model are summarized in Table 1. The computational implementation of the framework relies on several open-source libraries like FEniCS [48] for the finite element solution of the thermomechanical model, SOUPy [49] for the quadratic approximation of the design objective and implementation of chance constraint, and hIPPYLib [46] for the Gaussian random field, trace estimator and INCG optimizer [^1] The codes and links to the dependent libraries for this work can be found at: https://github.com/pce-lab/InsulationDesignUncertainty.
[…]
@@ 5.2 Quadratic and Monte Carlo approximations
Figure 5: Convergence plots corresponding to scenario in Figure 6 for the mean and variance of the design objective $Q$ : (a) standard Monte Carlo method (b) Monte Carlo with control variate and (c) Quadratic approximation. The uncertain parameter $m$ has the correlation length $L_{CR}=0.25$ and a variance of $\sigma^{2}={0.5}^{2}$.
@@
Figure 6: The top row shows the optimal design obtained for $\sigma=0.25$ for the three different approaches: (a) Sampled Average approximation (500 samples) (b) Monte Carlo correction (100 samples) and (c) Quadratic approximation ($N_{eig}^{Q}=25$). The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above. The uncertain parameter $m$ has a correlation length $L_{CR}=0.05$ for all three cases.
Figure 7: The top row shows the optimal design obtained for $\sigma=0.5$ for the three different approaches: (a) standard Monte Carlo using 500 samples, (b) Monte Carlo with control variate using 100 samples, and (c) Quadratic approximation using $N_{eig}^{Q}=25$. The bottom row shows the QoI distribution of the design objective $Q$ for the three different cases above. The uncertain parameter $m$ has a correlation length of $L_{CR}=0.25$ for all three cases.
To evaluate the accuracy and efficiency of the quadratic approximation and Monte Carlo estimations with a control variate in optimal design, we perform comparative analyses. Figure 5 shows that, as the sample size and number of dominant eigenvalues increase, both methods converge to similar mean and variance values, aligning with standard Monte Carlo estimations For large sample sizes, the estimated mean and variance from standard Monte Carlo (Figure 5(a)) and Monte Carlo with a control variate (Figure 5(b)) are nearly identical. However, the quadratic approximation as a control variate achieves lower error bounds with significantly fewer samples, as shown in the box plots. Notably, with only $N_{eig}^{Q}=25$ trace estimators, the quadratic approximation closely matches Monte Carlo results (Figure 5(c)), substantially reducing computational cost while maintaining accuracy.
Figures 6 and 7 compare the effects of different variances and correlation lengths of the uncertain parameter $m$ on the optimal design $d_{\text{opt}}$ indicating the spatial distribution of aerogel porosity and the probability distributions of the design objective $Q$ representing the thermal insulation performance of the thermal break. The optimal design is evaluated using standard Monte Carlo, Monte Carlo with control variate, and a quadratic Taylor approximation of the mean and variance of $Q$. In Figure 6 (low variance and short correlation length), the morphologies of $d_{\text{opt}}$ and probability distributions of $Q$ are nearly identical across all three methods. However, in Figure 7 (high variance and long correlation length), the quadratic approximation predicts different optimal design pattern with higher $Q$, while Monte Carlo with control variate yields results closely matching standard Monte Carlo but with significantly fewer samples (PDEs solved).
Interestingly, while the spatial pattern of the optimal design in Figure 6 appears nearly symmetric, the higher uncertainty in Figure 7 results in optimal designs that consistently position stronger material (lower aerogel porosity) along the horizontal axis, across all three approximation methods. This pattern reflects the anisotropy of the uncertain parameter, modeled with a higher correlation length along the layer deposition direction (see Figure 4). Also, the optimal designs at mean $\bar{m}=0$ are much similar in Figure 6 as compared to Figure 7 where there is a deviation in the design patterns, which shows that there is a trade-off favoring the reduction of the uncertainty over mean performance for our design formulation (7). These results demonstrate that the proposed optimization under uncertainty framework effectively captures spatial uncertainty, producing more reliable design solutions.
@@ 5.3 Scalability with respect to number of design parameters
We assess the scalability of the algorithm for optimization under uncertainty with respect to the design dimension, determined by the number of nodes in the spatial discretization of $d$, and governed by two primary factors: Firstly, the eigenvalue decay rate of the covariance-preconditioned Hessian, estimated via the randomized trace estimator, and secondly the convergence behavior of the INCG optimizer. Figure 8 illustrates the behavior of the chance constraint function in (9) with the number of steps in the continuation scheme, the gradient norm 53 and the absolute values of the generalized eigenvalues for the last step of chance convergence discussed in Algorithm 2. These metrics are presented for two design parameter dimensions, corresponding to the optimal design problem depicted in Figure 6 (c).
These results demonstrate that the quadratic approximation enables rapid decay of the critical chance, with the gradient norm converging below the tolerance threshold within 100 iterations for both mesh resolutions. The minimal dependence of convergence on mesh resolution highlights that the gradient-based INCG optimizer is effectively independent of the number of uncertain parameters. Additionally, Figure 8(c) shows a rapid eigenvalue decay, unaffected by the dimension of uncertain parameters, supporting the scalability of the randomized eigensolver used for trace approximation. Together, these dimension-independent factors suggest that leveraging quadratic Taylor approximations and stochastic optimization techniques enables a scalable approach for optimal design under high-dimensional uncertainty with the enforcement of chance constraints.
@@ 5.4 Optimal design: effect of chance constraint
Figure 9: Effect of the critical chance $\alpha_{c}$ on optimal design. The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of critical chance: (a) $\alpha_{c}=0.05$ (b) $\alpha_{c}=0.08$ and (c) $\alpha_{c}=0.1$. The bottom row represents the von Mises stress for the corresponding optimal designs in the top row. In all results $T_{cr}=22.5\>\text{MPa},L_{CR}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{RG}=1\times 10^{-5}$.
@@
Figure 10: Effect of the limiting critical stress $T_{cr}$ on optimal design. The top row shows the optimal design results obtained at the mean of the uncertain sample $\tilde{m}=0$ for different values of limiting critical stress: (a) $T_{cr}=20\>\text{MPa}$ (b) $T_{cr}=22.5\>\text{MPa}$ and (c) $T_{cr}=25\>\text{MPa}$. The bottom row represents the von Mises stress for the corresponding optimal designs in the top row. In all results $\alpha_{c}=0.05,L_{CR}=0.25,\sigma=0.25,\beta_{V}=0.1$ and $\beta_{RG}=1\times 10^{-5}$.
The chance constraint function defined in (9) incorporates parameters such as $\alpha_{c}$ and $T_{cr}$, which govern the enforcement of constrain to mitigate stress concentrations within the domain. In particular, the probability of chance $P(f(m,d)\geq 0)$ is limited by a threshold $\alpha_{c}$, where the function $f$ is the difference between the limiting critical stress $T_{cr}$ and the p-norm of the von Mises stress $T_{\text{pn}}$. To examine the influence of $\alpha_{c}$ on optimal design, a parametric study was conducted with constant parameters, as shown in Figure 9. The critical stress limit was set at $T_{cr}=22.5\>\text{MPa}$, while $\alpha_{c}$ was varied among 0.05, 0.08, and 0.1. This figure presents the optimal design, and von Mises stress distribution evaluated at the mean of the uncertain parameter $\bar{m}=0$. The results indicate that increasing critical chance $\alpha_{c}$ expands the regions of mechanically stronger material (highlighted in red), effectively reducing localized stresses across the component. However, this improvement in mechanical stability comes at the expense of insulation performance of the thermal break.
@@
The limiting critical stress, $T_{cr}$, defines the maximum permissible von Mises stress for the beam structure. Figure 10 shows the influence of $T_{cr}$ on the optimal designs, evaluated at the mean of the uncertain parameter $\bar{m}=0$, along with the corresponding von Mises stress distributions. The findings indicate that increasing the critical stress $T_{cr}$ results in the allocation of stronger insulating material (depicted in blue) across a larger region of the domain. This enhances the insulation performance but compromises the mechanical strength of the component. Thus, a higher value of $T_{cr}$ improves insulation efficiency at the expense of reduced mechanical stability. Figure 11 shows the states and porosity field evaluated at $\bar{m}=0$ and three samples for the scenario in Figure 10 (b). The states include the solid displacement magnitude $||\mathbf{u}_{s}||$, effective solid stress $||\mathbf{T}_{s}||$, solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$ and fluid pressure $p$.
@@ 5.5 Three-dimensional design of beam-insulator system
Figure 12: Evolution of design pattern with continuation scheme. Optimal design obtained at each step of the continuation scheme with increasing $\omega$ as shown in Figure 3. Random field as initial condition is shown in (a) and the results obtained for optimal design at mean $\tilde{m}=0$ for steps in adaptive optimization from step 1 to step 5 shown in (b)-(f). The uncertain parameter $m$ has a correlation length of $L_{CR}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$,the critical chance is $\alpha_{c}=0.1$, and the critical stress considered is $T_{cr}=12.5\>\text{MPa}$.
We demonstrate the efficiency and scalability of the proposed optimization under uncertainty framework by applying it to a three-dimensional design problem. Specifically, we consider the 3D thermal break scenario shown in Figure 1, with a thickness of 0.5. In this case, a uniform circular load is also applied to the middle of the top surface, simulating the influence of additional building envelope assemblies. The uncertain parameter $m$ is modeled with a horizontal correlation length of $L_{CR}=0.25$ and a variance of $\sigma^{2}={0.1}^{2}$. For the chance constraint, we set a critical chance of $\alpha_{c}=0.1$ and a critical stress of $T_{cr}=12.5\>\text{MPa}$. The finite element discretization of the 3D domain resulted in 919,464 design parameters.
@@
Figure 13: Optimal design patterns for the 3D scenario. Final design obtained after five steps of the continuation scheme, achieving convergence of the critical chance $\alpha_{c}$. The 3D domain’s finite element discretization, resulted in 919464 design parameters.
Figure 12 illustrates the optimal designs obtained at each step of the continuation scheme for a random field as initial condition, with increasing $\omega$ as shown in Figure 3. We tested different initial conditions and arrived at the same final pattern, indicating that for the scenario represented in Figure 12, we achieve global optima. Regions shown in red correspond to material with low porosity ($\phi_{f}=0.1$). As the continuation scheme progresses, the red regions start to form, increasing mechanical support in the domain and decreasing the p-norm of von Mises stress. Consequently, the probability in (10) approaches the critical chance threshold of $\alpha_{c}=0.1$. By step 5, the final design, shown in Figure 13, emerges, featuring a column-like structure of low-porosity material $\phi_{f}=0.1$, providing support aligned with the direction of the applied external load, and also along the left wall to provide support against the stress due to thermal expansion of the beam.
@@ 6 Conclusions
This study introduces an efficient and scalable computational framework for chance-constrained optimal design under high-dimensional uncertainty governed by finite element models. The framework optimizes the spatial distribution of a spatially correlated and uncertain porosity field in additively manufactured aerogel thermal breaks to maximize thermal insulation performance while mitigating stress concentrations. Two key strategies are implemented within the framework. The first utilizes a quadratic Taylor expansion of the design objective with respect to the uncertain parameters to efficiently estimate the mean and variance in the cost function, as well as the chance constraints. The second employs the quadratic approximation as a control variate to enhance the efficiency of Monte Carlo estimations. The framework incorporates a randomized algorithm to solve the generalized eigenvalue problem, enabling efficient computation of the trace of the covariance-preconditioned Hessian. Scalability with respect of high-dimensional uncertain parameter is achieved through the rapid decay of eigenvalues and the dimension-independent convergence of the optimization iterations.
@@
This study addresses the computational challenges of optimization under high-dimensional uncertainty by leveraging intrinsic dimensionality and incorporating chance constraints to prevent stress concentrations. The framework enhances optimal design reliability, offering a robust and scalable approach while paving the way for future advancements to refine and extend its capabilities.
Data Availability Statement
Some or all data, models, or code generated or used during the study are available in a repository online.